OpenStudy (anonymous):

Part A:The area of a square is (4x2 − 12x + 9) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. (5 points) Part B: The area of a rectangle is (16x2 − 9y2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. (5 points)

OpenStudy (anonymous):

@Haseeb96

OpenStudy (haseeb96):

4x^2 -12x +9 now 4 x 9 = 36 first take out the LCM of 36 it will be 2 x 2 x 3 x 3 4x^2 -6x -6x +9 2x(2x-3) -3(2x-3) (2x-3)(2x-3) it is showing that 2x-3 is the length for part A

OpenStudy (haseeb96):

for part B what you will do 16x^2 - 9y^2 you will use this formula here x^2 -y^2 = (x-y) (x+y) 16x^2 -9y^2 = (4x -3y) (4x +3y) this is the answer for part B

OpenStudy (haseeb96):

I hope you get what i explain here

OpenStudy (anonymous):

I do thank you so much can you help me with one more problem

OpenStudy (haseeb96):

okay

OpenStudy (anonymous):

An expression is shown below: f(x) = –16x2 + 22x + 3 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. (2 points) Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. (3 points) Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph. (5 points)

OpenStudy (haseeb96):

sorry idk about this one

OpenStudy (anonymous):

x intercept is when you set y=0

OpenStudy (anonymous):

okay but I am still confused on how you find part b and part c

OpenStudy (anonymous):